OpenStudy (anonymous):

PLEASEE HELPPPPPP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Question 3 (Multiple Choice Worth 2 points) Solve for x: the quantity of x plus 4 over 3 = 2 x = −2 x = 2 x = 2 over 3 x = -10 over 3

OpenStudy (amistre64):

posting the same question multiple times is considered to be spamish behaviour

OpenStudy (anonymous):

2~

OpenStudy (amistre64):

http://openstudy.com/code-of-conduct "Don't spam! Common forms of spam include: •Posting a question multiple times within a short period of time."

OpenStudy (anonymous):

I DIDNT MEAN TO I DID IT WRONG AT FIERST SO I HAD TO REPOST IR RIGHT WAY , OSMONDF ITS -2 ???

OpenStudy (amistre64):

they now provide you with an edit option to edit any mistakes :)

OpenStudy (amistre64):

x+4/3=2 is that x+4 all together?

OpenStudy (mertsj):

\[\frac{x+4}{3}=\frac{2}{1}\]

OpenStudy (mertsj):

Cross multiply and solve.

OpenStudy (mertsj):

\[x+4=6\]

OpenStudy (mertsj):

\[x=2\]

OpenStudy (anonymous):

-2 RIGHT ???

OpenStudy (amistre64):

the reson its considered spam is that it comes across as tho the person is trying to "overpower" other questions by having theirs on top.

OpenStudy (amistre64):

2/3 might be right depending on how the question is spose to be presented